OpenStudy (anonymous):

Part 1: Create a difference of squares binomial or a perfect square trinomial that can be factored multiple times. Part 2: Provide the factors of this polynomial. Part 3: Explain, in complete sentences, the process you used to create the binomial/trinomial.

OpenStudy (anonymous):

1. x^2 - 9 2. (x-3)(x+3) 3. Begin with two linear functions, f(x) = x - 3 and g(x) = x + 3. Then, multiply them by each other to get the expression (x-3)(x+3). Finally expand the expression to get x^2 - 9.

OpenStudy (anonymous):

Am I mistaken? Anyone?

OpenStudy (anonymous):

can it be factored multiple times?

OpenStudy (anonymous):

It can be factored n times. \[\tiny{\text{where n = 1}}\]

OpenStudy (anonymous):

Right. if you want one that can be factored multiple times, there's a way x^4 - 1 = (x^2 - 1)(x^2 + 1) = (x-1)(x+1)(x^2+1)

OpenStudy (anonymous):

the way to create that is by beginning with an ordinary difference of squares binomial like (x^2 - 1), and then substituting x by x^2 so that you get x^4 - 1

OpenStudy (anonymous):

give me one that can be factored multiple times, and do the part 1 and 2 and 3 to that.. cause it says multiple times

OpenStudy (anonymous):

I would start with the factors just like agdgdgdgwngo did. Maybe use (x+1)(x-1)(x+2)(x-2) and then factor it out to get the polynomial...

OpenStudy (anonymous):

lol so confusing !!!

OpenStudy (anonymous):

that's it; x^4 - 1 is a binomial that may be factored multiple times.